# Tight analyses of first-order methods with error feedback

### v1

@@ Appendix B Missing proofs
This section contains the proofs of Theorems 1 and 2. The proofs were obtained using by observing numerical results using the Lyapunov search procedure presented in Subsections A.2 and A.3. The resulting proofs are remarkably compact, and the main technical step consists of verifying an algebraic reformulation by hand. For ease of verification, we provide the reader with a script written in Wolfram Language [^3] Wolfram Language is the programming language used in Mathematica. The scripts can be used to verify our rates without a paid license using Wolfram Engine. that automatically performs those reformulations using a computer algebra system.
B.1 Proof of Theorem 1
@@
[…]
@@
Summing these inequalities with their multipliers, plugging in the update rules for $x_{k+1}$ and $e_{k+1}$, and using $\rho$ to denote the contraction factor we got in (9), we can rewrite the resulting inequality as:
$\rho\cdot\mathcal{V}(x_{k},e_{k})\geqslant\mathcal{V}(x_{k+1},e_{k+1})+a\cdot\|e_{k}-\frac{\rho-1}{a}(x_{k}-x_{\star})+\frac{2(\sqrt{\epsilon}-1)}{a(L+\mu)}g_{k}\|^{2},$ (46)
where
$a\coloneqq(\rho-\sqrt{\epsilon})\cdot\left(\frac{1+\sqrt{\epsilon}}{\sqrt{\epsilon}}\right).$ (47)
@@ B.2 Proof of Theorem 2
[…]
@@
Summing these inequalities with their multipliers, plugging in the update rules for $x_{k+1}$ and $d_{k+1}$, and using $\rho$ to denote the contraction factor we got in (12), we can rewrite the resulting inequality as:
$\rho\cdot\mathcal{V}(g_{k},d_{k})\geqslant\mathcal{V}(g_{k+1},d_{k+1})+a\cdot\|d_{k}+\frac{1}{a}((\epsilon+b)g_{k+1}-(\rho+b)g_{k})\|^{2},$ (60)
where
$b\coloneqq\frac{\lambda}{L-\mu}\cdot\left(\frac{1-\sqrt{\epsilon}}{1+\sqrt{\epsilon}}\right).$ (61)
and
$a\coloneqq\rho-\epsilon+2\eta\lambda(1-\sqrt{\epsilon})\frac{L\mu}{L-\mu}.$ (62)
@@
Appendix C Additional numerical results
@@
  • Subsection C.1 presents additional performance plots for all methods and explains how the plots in the main paper were generated.
  • Subsection C.2 provides illustrations demonstrating that our Lyapunov functions remain tight over multiple iterations.
  • Subsection C.3 includes additional tables further confirming the tightness of our Lyapunov functions.
  • Subsection C.4 shows plots illustrating the optimality of our step size.
@@ C.1 Performance plots
To identify the area of non-convergence in the plots, we check whether a cycle exists for each pair of $\eta$ and $\epsilon$. This is done by following the procedure outlined in Goujaud et al. [63]: we compute the worst-case performance of the metric $-\|x_{k}-x_{0}\|^{2}$ for CGD, $-\|x_{k}-x_{0}\|^{2}-\|e_{k}-e_{0}\|^{2}$ for $\mathrm{EF}$, and $-\|x_{k}-x_{0}\|^{2}-\|d_{k}-d_{0}\|^{2}$ for $\mathrm{EF^{21}}$. If this value falls below a threshold (set to $10^{-3}$), we conclude that a cycle is present. In our experiments, cycles of length 2 were successfully identified for all methods, and these matched precisely with the regions of the contour plots where $\rho>1$.
@@
C.2 Multi-step Lyapunov analysis
In this section, we show that our simple Lyapunov functions achieve the claimed convergence rate over multiple iterations. Specifically, we use PEPit to compute the contraction factor achieved by the Lyapunov function after $k$ iterations and compare it to the theoretical rate $\rho^{k}$, where $\rho$ is the single-step contraction factor. The exact match between these quantities in Figures 6 and 7 confirms that our single-step analysis accurately characterizes the worst-case performance over multiple iterations on these Lyapunov functions.
@@ C.2 Multi-step Lyapunov analysis
C.3 Lyapunov function class tightness
@@
The tables report the maximum absolute difference in contraction factors between our Lyapunov function and the optimal one, over a range of $\epsilon$ and $\eta$ values specified in the captions. All contraction factors were computed using PEPit, and the procedure for the uncontrained Lyapunov functions is the one outlined in Subsection C.1. Points where either Lyapunov function yields a contraction factor greater than 1 were excluded from the computation of the maximum absolute difference.
$\kappa=2.0$ $\kappa=4.0$ $\kappa=10.0$
@@
Table 3: Maximum absolute difference in contraction factor for $\mathrm{EF}$ when comparing the general Lyapunov function—constructed using any combination of state terms specified in Subsection 3.1—to the simplified Lyapunov function defined in Theorem 1. The results are computed over a line with $\epsilon\in[0.01,0.99]$ and with $\eta$ set to the optimal step size for $L=1$, and $\mu=0.5,0.25,0.1$.
$\kappa=2.0$ $\kappa=4.0$ $\kappa=10.0$
@@
Table 4: Maximum absolute difference in contraction factor for $\mathrm{EF^{21}}$ when comparing the general Lyapunov function—constructed using any combination of state terms specified in Subsection 3.1—to the simplified Lyapunov function defined in Theorem 2. The results are computed over a line with $\epsilon\in[0.01,0.99]$ and with $\eta$ set to the optimal step size for $L=1$, and $\mu=0.5,0.25,0.1$.
C.4 Step size comparison
In this section, we compare the theoretically optimal step sizes we propose for our methods with empirically optimal step sizes determined through numerical experiments. To compute the empirical optima, we evaluate a grid of $\eta$ and $\epsilon$ values and select the step size that minimizes the contraction factor achieved by our simplified Lyapunov functions. The results of that experiment are found in Figure 8.

### v2

@@ Appendix C Missing proofs
This section contains the proofs of Theorems˜1 and 2. The proofs were obtained using by observing numerical results using the Lyapunov search procedure presented in Subsections˜A.2 and A.3. The resulting proofs are remarkably compact, and the main technical step consists of verifying an algebraic reformulation by hand. For ease of verification, we provide the reader with scripts written in Wolfram Language [^4] Wolfram Language is the programming language used in Mathematica. The scripts can be used to verify our rates without a paid license using Wolfram Engine. that automatically performs those reformulations using a computer algebra system.
C.1 Proof of Theorem 1
@@
$\begin{aligned} I_{\mathcal{F}_{\mu,L}}^{(1)}\coloneqq f(x_{k})-f_{\star}-\nabla f(x_{k})^{\top}(x_{k}-x_{\star})+\frac{1}{2L}\|\nabla f(x_{k})\|^{2}\\ +\frac{\mu}{2(1-\mu/L)}\|x_{k}-x_{\star}-\frac{1}{L}\nabla f(x_{k})\|^{2}\leqslant 0,\end{aligned}$ $:\lambda$
@@
Certified using WolframScriptCertified using WolframScriptCAS  2 CAS  2 Summing these inequalities with their multipliers, plugging in the update rules for $x_{k+1}$ and $e_{k+1}$, and using $\rho$ to denote the contraction factor we got in (9), we can rewrite the resulting inequality as: $\rho\cdot\mathcal{V}(x_{k},e_{k})\geqslant\mathcal{V}(x_{k+1},e_{k+1})+a\cdot\|e_{k}-\frac{\rho-1}{a}(x_{k}-x_{\star})+\frac{2(\sqrt{\epsilon}-1)}{a(L+\mu)}g_{k}\|^{2},$ (51) where $a\coloneqq(\rho-\sqrt{\epsilon})\cdot\left(\frac{1+\sqrt{\epsilon}}{\sqrt{\epsilon}}\right).$ (52)
@@ C.2 Proof of Theorem 2
$\begin{aligned} I_{\mathcal{F}_{\mu,L}}^{(1)}\coloneqq f(x_{k})-f(x_{k+1})+\frac{\|\nabla f(x_{k+1})-\nabla f(x_{k})\|^{2}}{2L}+\nabla f(x_{k})^{\top}(x_{x+1}-x_{k})\\ +\frac{\mu}{2(1-\mu/L)}\|x_{k}-x_{k+1}-\frac{1}{L}(\nabla f(x_{k})-\nabla f(x_{k+1}))\|^{2}\leqslant 0,\end{aligned}$ $:\lambda^{\prime}$
$\begin{aligned} I_{\mathcal{F}_{\mu,L}}^{(2)}\coloneqq f(x_{k+1})-f(x_{k})+\frac{\|\nabla f(x_{k})-\nabla f(x_{k+1})\|^{2}}{2L}+\nabla f(x_{k+1})^{\top}(x_{k}-x_{k+1})\\ +\frac{\mu}{2(1-\mu/L)}\|x_{k+1}-x_{k}-\frac{1}{L}(\nabla f(x_{k+1})-\nabla f(x_{k}))\|^{2}\leqslant 0,\end{aligned}$ $:\lambda^{\prime}$
@@
Certified using WolframScriptCertified using WolframScriptCAS  2 CAS  2 Summing these inequalities with their multipliers, plugging in the update rules for $x_{k+1}$ and $d_{k+1}$, and using $\rho$ to denote the contraction factor we got in (12), we can rewrite the resulting inequality as: $\rho\cdot\mathcal{V}(g_{k},d_{k})\geqslant\mathcal{V}(g_{k+1},d_{k+1})+a\cdot\|d_{k}+\frac{1}{a}((\epsilon+b)g_{k+1}-(\rho+b)g_{k})\|^{2},$ (65) where $b\coloneqq\frac{\lambda}{L-\mu}\cdot\left(\frac{1-\sqrt{\epsilon}}{1+\sqrt{\epsilon}}\right).$ (66) and $a\coloneqq\rho-\epsilon+2\eta\lambda(1-\sqrt{\epsilon})\frac{L\mu}{L-\mu}.$ (67)
@@
C.3 Extension to stochastic compressors
Note that all the results of Theorems˜1 and 2 were proven using the deterministic version of Assumption˜1. Any deterministic compressor satisfies that assumption, and our lower bounds thus remain valid in this case. To show that the given rates are tight, we have to replace the inequalities $I_{\mathcal{C}}$ from the respective proofs with their stochastic counterparts:
$I_{\mathcal{C}}^{\mathrm{EF}}\coloneqq\mathbb{E}\left[\|e_{k+1}\|^{2}\right]-\epsilon\|e_{k}+\eta\nabla f(x_{k})\|^{2}\leqslant 0,$
$I_{\mathcal{C}}^{\mathrm{EF^{21}}}\coloneqq\mathbb{E}\left[\|\nabla f(x_{k+1})-d_{k}-\mathcal{C}(\nabla f(x_{k+1})-d_{k})\|^{2}\right]-\epsilon\|\nabla f(x_{k+1})-d_{k}\|^{2}\leqslant 0.$
Certified using WolframScriptCertified using WolframScriptCAS  2 CAS  2 For $\mathrm{EF}$, we need to show the following: $\rho\cdot\mathbb{E}\mathcal{V}(x_{k},e_{k})\geqslant\mathbb{E}\mathcal{V}(x_{k+1},e_{k+1}),$
Certified using WolframScriptCertified using WolframScriptCAS  2 CAS  2 And for $\mathrm{EF^{21}}$ we need to show the following: $\rho\cdot\mathbb{E}\mathcal{V}(g_{k},d_{k})\geqslant\mathbb{E}\mathcal{V}(g_{k+1},d_{k+1}).$
The same exact proof as used in the deterministic case holds for both due to the linearity of expectation. Wolfram Language scripts verifying this fact are available in the source code repository for this article.
Appendix D Additional numerical results
@@
  • Subsection˜D.1 presents additional performance plots for all methods and explains how the plots in the main paper were generated.
  • Subsection˜D.2 provides illustrations demonstrating that our Lyapunov functions remain tight over multiple iterations.
  • Subsection˜D.3 includes additional tables further confirming the tightness of our Lyapunov functions.
  • Subsection˜D.4 shows plots illustrating the optimality of our step size.
  • Subsection˜D.5 numerically demonstrates that the type of Lyapunov function used in Richtárik et al. [14] achieves a worse rate of convergence than the one used in this work.
@@ D.1 Performance plots
To identify the area of non-convergence in the plots, we check whether a cycle exists for each pair of $\eta$ and $\epsilon$. This is done by following the procedure outlined in Goujaud et al. [66]: we compute the worst-case performance of the metric $\|x_{k}-x_{0}\|^{2}$ for CGD, $\|x_{k}-x_{0}\|^{2}-\|e_{k}-e_{0}\|^{2}$ for $\mathrm{EF}$, and $\|x_{k}-x_{0}\|^{2}-\|d_{k}-d_{0}\|^{2}$ for $\mathrm{EF^{21}}$. If this value falls below a threshold (set to $10^{-3}$), we conclude that a cycle is present. In our experiments, cycles of length 2 were successfully identified for all methods, and these matched precisely with the regions of the contour plots where $\rho>1$.
@@
D.2 Multi-step Lyapunov analysis
In this section, we show that our simple Lyapunov functions achieve the claimed convergence rate over multiple iterations. Specifically, we use PEPit to compute the contraction factor achieved by the Lyapunov function after $k$ iterations and compare it to the theoretical rate $\rho^{k}$, where $\rho$ is the single-step contraction factor. The exact match between these quantities in Figures˜9 and 10 confirms that our single-step analysis accurately characterizes the worst-case performance over multiple iterations on these Lyapunov functions.
@@ D.2 Multi-step Lyapunov analysis
D.3 Lyapunov function class tightness
@@
The tables report the maximum absolute difference in contraction factors between our Lyapunov function and the optimal one, over a range of $\epsilon$ and $\eta$ values specified in the captions. All contraction factors were computed using PEPit, and the procedure for the uncontrained Lyapunov functions is the one outlined in Subsection˜D.1. Points where either Lyapunov function yields a contraction factor greater than 1 were excluded from the computation of the maximum absolute difference.
$\kappa=2$ $\kappa=4$ $\kappa=10$
@@
Table 3: Maximum absolute difference in contraction factor for $\mathrm{EF}$ when comparing the general Lyapunov function—constructed using any combination of state terms specified in Subsection˜3.1—to the simplified Lyapunov function defined in Theorem 1. The results are computed over a line with $\epsilon\in[0.01,0.99]$ and with $\eta$ set to the optimal step size for $L=1$, and $\mu=0.5,0.25,0.1$.
$\kappa=2$ $\kappa=4$ $\kappa=10$
@@
Table 4: Maximum absolute difference in contraction factor for $\mathrm{EF^{21}}$ when comparing the general Lyapunov function—constructed using any combination of state terms specified in Subsection˜3.1—to the simplified Lyapunov function defined in Theorem 2. The results are computed over a line with $\epsilon\in[0.01,0.99]$ and with $\eta$ set to the optimal step size for $L=1$, and $\mu=0.5,0.25,0.1$.
D.4 Step size comparison
In this section, we compare the theoretically optimal step sizes we propose for our methods with empirically optimal step sizes determined through numerical experiments. To compute the empirical optima, we evaluate a grid of $\eta$ and $\epsilon$ values and select the step size that minimizes the contraction factor achieved by our simplified Lyapunov functions. The results of that experiment are found in Figure˜11.